Okay this is a function of a right triangle where side a is equal to 16 and its corresponding angle is 45 degrees side b and the hypoteneous are unknown. Is there a way to solve for the two missing sides?
Answer by fractalier(6550)

Sure. If it is a right triangle and one angle is 45 degrees, the other must also be 45 degrees (all angles in a triangle add to 180, right?)...
Now if the base angles are equal, the sides must also be...we have an isosceles right triangle...
So side b is also 16.
We can use the Pythagorean Theorem to find the hypotenuse...
a^2 + b^2 = c^2
16^2 + 16^2 = c^2
c^2 = 512
c = sqrt(512) = 16sqrt(2)